### v2

# Conflicts in Texts: Data, Implications and Challenges

Abstract
As NLP models become increasingly integrated into real-world applications, it becomes clear that there is a need to address the fact that models often rely on and generate conflicting information. Conflicts could reflect the complexity of situations, changes that need to be explained and dealt with, difficulties in data annotation, and mistakes in generated outputs. In all cases, disregarding the conflicts in data could result in undesired behaviors of models and undermine NLP models’ reliability and trustworthiness. This survey categorizes these conflicts into three key areas: (1) natural texts on the web, where factual inconsistencies, subjective biases, and multiple perspectives introduce contradictions; (2) human-annotated data, where annotator disagreements, mistakes, and societal biases impact model training; and (3) model interactions, where hallucinations and knowledge conflicts emerge during deployment. While prior work has addressed some of these conflicts in isolation, we unify them under the broader concept of conflicting information, analyze their implications, and discuss mitigation strategies. We highlight key challenges for developing conflict-aware and robust NLP systems, and propose concrete research directions to address them.
1 Introduction
The rapid advancement of natural language processing (NLP), particularly with the rise of large language models (LLMs), has led to their widespread adoption in daily tasks, information retrieval, and decision-making processes. However, the increasing complexity of these models reveals various types of conflicts at multiple stages, including training, annotation, and model interaction, affecting the reliability and trustworthiness of downstream applications. For example, training models on data containing factual contradictions, annotation disagreements, or prompts that contradict a model’s parametric knowledge can introduce inconsistencies with unpredictable consequences Pavlick and Kwiatkowski (2019); Sap et al. (2019).
Figure 1: Examples of the three different areas of conflicts discussed in this work. The first example describes a case where two different entities of the same name are found naturally on the web, the second example elaborates the annotation disagreement in a sentiment analysis task, and the third showcases a knowledge conflict between the context and memory of LLMs during model interactions.
Existing work on conflicts in NLP tends to focus on specific issues, such as annotation disagreements Uma et al. (2021); Klie et al. (2023), hallucinations and factuality Zhang et al. (2023); Wang et al. (2023), and knowledge conflicts Xu et al. (2024); Feng et al. (2024), without synthesizing these problems into a broader perspective. In this survey, we conceptualize these diverse challenges under the umbrella of conflicting information and analyze their origins, implications, and mitigation strategies.
{forest}
t
Figure 2: Taxonomy of conflicts in texts.
To ensure comprehensive and representative coverage of conflicts in NLP, we first establish a high-level categorization encompassing three primary sources: (1) natural conflicts present in web data, (2) conflicts arising from human annotation, and (3) conflicts emerging from model interactions. Notably, conflicts found in natural web texts and human-annotated datasets are primarily present in the training data—i.e., the inputs to models—whereas conflicts involving model interactions can arise in various forms, such as inconsistencies between model outputs and their inputs, contradictions among multiple outputs, or conflicts within the outputs themselves. For each category, we identify influential and widely cited survey papers as initial seed works  Uma et al. (2021); Klie et al. (2023); Zhang et al. (2023); Xu et al. (2024); Feng et al. (2024); Wang et al. (2023). Building upon these seeds, we systematically trace and incorporate the most impactful and representative studies for each type of conflict through citation chaining and targeted literature searches across major databases. This approach enables us to synthesize developments in each category and connect them, thereby providing an integrated discussion of current challenges, impacts on downstream tasks, and promising future directions for conflict-aware AI systems (See Appendix B for more discussion about our survey methodology).
The abundance of online data is accompanied by inherent conflicts, stemming from diverse sources, interpretations, and biases. These conflicts manifest as factual conflicts, such as semantic ambiguities Pavlick and Tetreault (2016); Min et al. (2020) [^1] The previous version on ACL Anthology has a wrong paper title for a reference here. and factual inconsistencies Pham et al. (2024); Liu et al. (2024), or as conflicts in opinions related to political ideologies Entman (1993); Recasens and et al. (2013) and perspectives Chen et al. (2019); Liu et al. (2021). Factual conflicts are particularly prevalent in open-domain question answering (QA) and retrieval-augmented generation (RAG) systems Chen et al. (2017), where aggregating knowledge from multiple sources introduces inconsistencies Liu et al. (2024). These challenges highlight the need for conflict-aware retrieval and reasoning mechanisms to improve model reliability Xie et al. (2024). Unlike factual conflicts, opinionated disagreements reflect the variability in human interpretation, beliefs, and ideological stances Chen et al. (2019); Fan et al. (2019). The presence of conflicting viewpoints complicates tasks such as summarization, sentiment analysis, and dialogue generation, where maintaining coherence and neutrality is crucial Liu et al. (2021); Lee et al. (2022). Furthermore, the uneven distribution and biases of web data also affects models to behave from a Western perspective Ramaswamy et al. (2023); Mihalcea et al. (2025).
Another significant conflict arises in human-annotated data. For instance, annotation disagreements persists in both subjective and seemingly objective NLP tasks Mostafazadeh Davani et al. (2022). Disagreements are widespread in sentiment analysis Wan et al. (2023), hate speech detection Sap et al. (2022), and even natural language inference (NLI) Pavlick and Kwiatkowski (2019). Models trained on aggregated (e.g. majority-vote) labels struggle with ambiguous or high-disagreement examples, often treating them as hard-to-learn or mislabeled Anand et al. (2024). Pavlick and Kwiatkowski (2019) also find that standard NLI models’ uncertainty does not reflect the true ambiguity present in human opinions, leading to overconfidence in contentious cases. In addition, annotation biases—such as those related to race, gender, and geography—skew model predictions and reinforce societal biases Buolamwini and Gebru (2018); Sap et al. (2022); Pei and Jurgens (2023). These issues highlight the need for fair and representative annotations that capture the complexity of human disagreement.
Conflicts also emerge during interactions with models, manifesting as knowledge conflicts between model memories and contexts, and hallucinations in generated outputs. Knowledge conflicts arise when a model’s internal memory contradicts external contextual evidence, as shown by Longpre et al. (2021), who found that models often overly depend on memorized knowledge, leading to hallucinations. Neeman et al. (2023) proposed separating parametric and contextual knowledge to improve interpretability, while Xie et al. (2024) examined LLMs’ confirmation bias, showing how models inconsistently handle contradictory evidence. Additionally, hallucinations—ranging from factual inconsistencies Lin et al. (2022); Ouyang et al. (2022) to contextual hallucinations Maynez et al. (2020); Kryscinski et al. (2020)—further undermine model reliability. Various mitigation strategies have been proposed, including retrieval augmentation Lewis et al. (2020); Shuster et al. (2021), hallucination detection Manakul et al. (2023), and knowledge graph-based verification Guan et al. (2024).
In this survey, we systematically examine the landscape of conflicts in NLP by categorizing them into three primary sources. For each conflict type, we detail how such conflicts arise and in what forms they take (origins), the challenges they pose (implications), and the strategies developed to address them (mitigation). We present a comprehensive taxonomy in Figure 2, as well as structured summary tables—Table 1, Table 2, and Table 3—that synthesize datasets, methodologies, and analysis from prior work. By offering a unified framework for understanding and addressing conflicting information in NLP, this survey contributes to the development of conflict-aware frameworks for data collection, model training, and model usage, ultimately enhancing the fairness and reliability of NLP.
2 Conflicts in Natural Texts on the Web
Conflicts in natural texts on the web manifest in diverse ways, reflecting the inherent complexity and subjectivity of human language. They can broadly be categorized into factual conflicts, which revolve around factual discrepancies caused by various reasons, and conflicts in opinions, which pertain to divergent perspectives or biases.
2.1 Factual Conflicts
2.1.1 Origins
Ambiguity
Ambiguity is a root cause of factual conflict. When a query or piece of data lacks clarity about entities or context, a model can produce conflicting answers. A clear demonstration of how ambiguity induces conflicts is context dependence. For example, an ambiguous question of "which COVID-19 vaccine was the first to be authorized by our government?" can have conflicting answers depending on different geographical contexts Zhang and Choi (2021).
Min et al. (2020) was the first work to study the effects of ambiguity in open domain question answering. They introduced AmbigQA, a dataset highlighting that over half of the open-domain, natural questions are ambiguous, with diverse sources of ambiguity such as event and entity references. Zhang and Choi (2021) proposed the SituatedQA task, showing that a significant fraction of open-domain questions are valid only under particular temporal or geographic contexts. Many other work specifically focus on the temporal aspect of ambiguity, benchmarking and evaluating models’ awareness and adaptation to time-sensitive questions Chen et al. (2021); Liska et al. (2022); Kasai et al. (2023).
Contradictory Evidence
Conflicts in NLP systems arise when information on the web presents conflicting evidence towards a factual question. This issue is particularly prevalent in open-domain question answering settings, where models must navigate inconsistencies across diverse information sources. For example, Liu et al. (2024) find that 25% of unambiguous factual questions queried on Google retrieve conflicting evidence from multiple sources.
Researchers have proposed different datasets to systematically study how NLP models handle such conflicts. Li et al. (2024b) introduce ContraDoc, a human-annotated dataset of long documents with internal contradictions; Pham et al. (2024) propose WhoQA, a benchmark dataset that constructs conflicts by formulating questions about a shared property among entities with the same name (e.g. "Who is George Washington?"); and Liu et al. (2024) construct QACC, a human-annotated dataset of conflicting results retrieved by Google. Beyond empirical datasets, several studies have proposed synthetic approaches to simulate conflicts through entity substitution Chen et al. (2022a); Hong et al. (2024), machine-generated conflicting evidence Pan et al. (2023); Wan et al. (2024a); Hong et al. (2024), and pre-defined rule-based templates Kazemi et al. (2023).
2.1.2 Implications and Mitigation
Implications
Factual conflicts pose significant challenges for NLP systems. Pre-trained language models accurately detect context-dependent questions but fall short when answering queries requiring temporal context, performing notably below human levels Zhang and Choi (2021). Additionally, large language models (LLMs) often exhibit confirmation bias, favoring retrieved information that aligns with their parametric memory despite contradictory evidence Xie et al. (2024). Consequently, conflicting information sources severely impact retrieval-augmented generation (RAG) frameworks, significantly degrading model performance even with minimal misinformation exposure Pham et al. (2024); Liu et al. (2024); Li et al. (2024b); Pan et al. (2023).
Mitigation
To address these challenges, various mitigation strategies have been proposed. Effective methods include fine-tuning calibrators for selective abstention Chen et al. (2022a), employing a "disambiguate-then-answer" pipeline to detect ambiguity proactively Cole et al. (2023), and developing time-aware models that condition responses on timestamps to manage outdated information Dhingra et al. (2022). Further robustness improvements have been achieved through fine-tuning discriminators or prompting GPT-3.5 models to explicitly recognize conflicting evidence Hong et al. (2024), as well as incorporating human-written explanations in fine-tuning processes to enhance models’ reasoning capabilities Liu et al. (2024).
2.2 Conflicts in Opinions
2.2.1 Origins
Perspectives
Individuals and communities often hold diverse perspectives on the same issue. Such diversity is evident in online discussions and debates, where the multiplicity of viewpoints can lead to conflicting opinions. For instance, on controversial topics such as "Animals should have lawful rights," people express varying stances Chen et al. (2019), posing challenges for downstream tasks like summarization where consolidating viewpoints and presenting unbiased information are crucial Liu et al. (2021); Lee et al. (2022).
Several studies have explored perspectives in the context of conflicting information. Chen et al. (2019) introduce the task of substantiated perspective discovery, where systems identify diverse, evidence-supported stances on a claim, and release the PERSPECTRUM dataset using online debates and search results. Wan et al. (2024a) propose ConflictingQA, a dataset of controversial questions paired with real-world documents that present divergent facts, arguments, and conclusions. Plepi et al. (2024) examine perspective-taking in contentious online discourse, curating a corpus of 95k conflict scenarios annotated with users’ self-reported backgrounds. Liu et al. (2021) present MultiOpEd, a corpus of 1,397 controversial topics, each paired with opposing editorials and concise summaries capturing their core perspectives.
Framing Bias
A specific example of how differing opinions are conveyed and expanded is framing bias, a mechanism in which news media shape interpretations by emphasizing certain aspects of information over others Entman (1993). In a polarized media environment, partisan media outlets deliberately frame news stories in a way to advance certain political ideologies Jamieson et al. (2007); Levendusky (2013); Liu et al. (2019).
Numerous studies have investigated different aspects of media bias. Card et al. (2015) introduce the Media Frames Corpus (MFC), a collection of news articles annotated with 15 general-purpose framing dimensions across three policy issues, enabling computational analysis of media framing. Liu et al. (2019) present the Gun Violence Frame Corpus (GVFC), a dataset of news headlines annotated by domain experts to capture framing in gun violence reporting. Fan et al. (2019) examine informational bias—bias conveyed through content selection and structure—and release BASIL, a dataset of 300 news articles annotated with 1,727 bias spans, demonstrating that informational bias is more prevalent than lexical bias.
2.2.2 Implications and Mitigation
Implications
Analysis of PERSPECTRUM reveals significant natural language understanding challenges, as human performance substantially outperforms machine baselines at identifying diverse, evidence-supported perspectives Chen et al. (2019). Furthermore, when selecting real-world evidence for controversial questions, LLMs predominantly prioritize the relevance of the evidence to the query, often disregarding stylistic attributes such as the presence of scientific references or a neutral tone Wan et al. (2024a). In addition, the distribution and biases of web data also affects models to behave from a Western perspective Ramaswamy et al. (2023); Mihalcea et al. (2025). Studies have shown that LLMs’ outputs skew toward the values of Western English-speaking countries Tao et al. (2024); Naous et al. (2024), and misalignment is more pronounced for underrepresented personas and on culturally sensitive topics such as social values Al Kuwatly et al. (2020). Furthermore, LLMs often provide inconsistent answers to the same question when prompted in different languages Li et al. (2024a); AlKhamissi et al. (2024); Eloundou et al. (2025), revealing conflicting cultural perspectives within a single model.
Mitigation
Several studies have proposed methods to address conflicts in perspectives and ideological bias. Liu et al. (2021) show that auxiliary tasks improve perspective summarization quality, while Chen et al. (2022b) propose a retrieval paradigm that clusters documents by viewpoint, revealing users’ preference for diverse perspectives over relevance-ranked lists. Jiang et al. (2023b) generate opinion summaries by selecting review subsets based on sentiment polarity and contrast, producing balanced pros, cons, and verdicts. Plepi et al. (2024) demonstrate that conditioning generation on users’ personal contexts yields more empathetic and appropriate responses than general-purpose models.
To mitigate framing and ideological bias, Milbauer et al. (2021) uncover nuanced worldview differences across communities by identifying multiple axes of polarization beyond the traditional left–right spectrum. Liu et al. (2022b) pre-train models for ideology detection by comparing reporting on the same events across partisan sources. Chen et al. (2023) disentangle content from style to enable ideology classification under data scarcity and bias. Lee et al. (2022) employ hierarchical multi-task learning to neutralize bias from news titles to article bodies, while Liu et al. (2023) construct neutral event graphs by synthesizing perspectives across ideological divides.
3 Conflicts in Human-Annotated Texts
Conflicts in human-annotated texts largely arise from two sources: annotation disagreements and societal or ethical biases. Disagreements stem from linguistic ambiguity, annotator backgrounds, and task design, while biases reflect systematic demographic or ideological influences that can skew labeling in consistent ways. Though conceptually distinct, these sources often interact—biases may amplify disagreement or entrench disparities. Differentiating between them is essential for understanding annotation-related conflicts and for developing more reliable and equitable NLP datasets.
3.1 Origins
Annotation Disagreement
The subjective nature of human judgment introduces variability and disagreement into annotated data Kahneman (2021). In NLP, such disagreements arise from linguistic ambiguity, annotator backgrounds, task design, and dataset curation practices. Uma et al. (2021) survey disagreements across NLP and vision tasks, identifying subjective ambiguity and annotator diversity as key contributors. Sandri et al. (2023) classify disagreements in offensive language detection as stemming from inherent ambiguity, annotation errors, or contextual gaps, highlighting that some disagreements reflect hard-to-classify content, while others indicate correctable issues. Similarly, Jiang and de Marneffe (2022) categorize NLI disagreements into linguistic uncertainty, annotator bias, and task design, showing that much of the observed noise is systematic and predictable.
Task formulation also plays a critical role. Dsouza and Kovatchev (2025) find that label disagreement in reinforcement learning from human feedback (RLHF) is shaped by annotator selection and task phrasing. Demographic and ideological factors further influence disagreements. Pavlick and Kwiatkowski (2019) argue that many NLI disagreements reflect genuine linguistic ambiguity and individual variation rather than annotation error. Sap et al. (2022) demonstrate that annotators’ personal beliefs and identities affect toxicity judgments, while Wan et al. (2023) show that demographic features significantly improve disagreement prediction.
Ethical and Societal Biases
Human-annotated texts also encode societal biases related to race, gender, and geography, which can significantly skew model predictions and downstream decisions Buolamwini and Gebru (2018). Sap et al. (2022) show that annotators’ ideological and racial identities influence toxicity judgments, with conservative annotators less likely to flag anti-Black slurs and more likely to misclassify African American English (AAE) as offensive. Thorn Jakobsen et al. (2022) examine how annotation guidelines interact with annotator demographics, demonstrating that even well-designed tasks can elicit systematically different responses across groups, highlighting the need for inclusive task design. Pei and Jurgens (2023) introduce POPQUORN, a dataset designed to assess demographic effects on annotation across NLP tasks, and find that annotator attributes—such as age, gender, race, and education—account for substantial variance in labeling behavior.
3.1.1 Implications and Mitigation
Implications
Early research has underscored the impact of annotator disagreement on data quality and model performance Artstein and Poesio (2008); Pustejovsky and Stubbs (2012); Plank et al. (2014). Pavlick and Kwiatkowski (2019) show that standard NLI models fail to capture the true uncertainty present in human judgments, leading to overconfidence on contentious examples. Similarly, Anand et al. (2024) find that models trained on single “gold” labels perform poorly and exhibit lower confidence on high-disagreement instances, often treating them as mislabeled or hard to learn. Sap et al. (2019) demonstrate how annotator bias can yield discriminatory outcomes: tweets in African American English (AAE) are frequently misclassified as toxic, a bias inherited by models that disproportionately flag content from Black authors. Additionally, many widely used NLP datasets exhibit a strong Western-centric skew Faisal et al. (2022), causing models to generalize poorly to underrepresented regions—for example, excelling on questions about New York or London, but failing on Nairobi or Manila due to lack of exposure.
Mitigation
Prior work has explored collecting multiple labels per data item to capture annotation variability and improve data quality. Probabilistic models have been developed to infer true labels by accounting for annotator expertise and label noise Sheng et al. (2008). Mostafazadeh Davani et al. (2022) propose a multi-task neural network that models each annotator’s labels individually while sharing a common representation, preserving disagreement in training. Similarly, studies show that models trained on soft labels—i.e., full label distributions reflecting annotator disagreement—consistently outperform those trained on aggregated labels Uma et al. (2021); Fornaciari et al. (2021).
4 Conflicts during Model Interactions
Conflicts during model interactions primarily manifest as knowledge conflicts and hallucinations, each posing distinct challenges. Knowledge conflicts occur when a model’s parametric memory contradicts contextual input or when inconsistencies arise across models, whereas hallucinations occur when outputs deviate from real-world facts or the given input. Differentiating these two types of conflict clarifies their underlying causes and helps guide targeted mitigation strategies.
4.1 Knowledge Conflicts
4.1.1 Origins
Context vs. Memory
A common type of knowledge conflict arises when a model’s prompt (contextual knowledge) contradicts what the model has learned and stored in its parameters (parametric knowledge) Longpre et al. (2021); Chen et al. (2022a). One prevalent cause of such conflicts is the presence of updated information Chen et al. (2021); Lazaridou et al. (2021); Luu et al. (2022), where newly available knowledge contradicts models’ previously learned knowledge.
Recent studies have developed many evaluation frameworks and datasets to assess LLMs’ behaviors in this scenario through different methods, including entity substitution Longpre et al. (2021); Chen et al. (2022a); Wang et al. (2024), adversarial perturbation Chen et al. (2022a); Xie et al. (2024), misinformation injection Pan et al. (2023), and machine generation Qian et al. (2024); Ying et al. (2024); Tan et al. (2024).
Within and Across Models
Conflicts may also arise across or within model knowledge bases. Cohen et al. (2023) explore how different LLMs encode different knowledge and can be used to fact-check one another, uncovering inconsistencies indicative of factual errors. Zhu et al. (2024) examine cross-modality conflicts in vision-language models, attributing discrepancies between visual and textual components to separate training regimes and distinct data sources. Even within a single model, contradictions can emerge: Zhao et al. (2024) detect intra-model inconsistencies by paraphrasing queries and observing divergent answers across prompts.
4.1.2 Implications and Mitigation
Implications
Interestingly, different studies of knowledge conflicts present seemingly contradictory findings. Some studies claim that models often excessively rely on parametric memory when observing conflicts with contextual knowledge Longpre et al. (2021); Some other studies posit that LLMs tend to ground their answers in retrieved documents in this scenario Chen et al. (2022a); Qian et al. (2024); Tan et al. (2024); or even both – LLMs are highly receptive to context when it is the only evidence presented in a coherent way, but also demonstrate a strong confirmation bias toward parametric memory when both supportive and contradictory evidence to their parametric memory are present Xie et al. (2024).
Mitigation
Several approaches have been proposed to mitigate the impact of knowledge conflicts. Longpre et al. (2021) reduce memorization by augmenting training data through corpus substitution. Chen et al. (2022a) introduce a calibrator that abstains from prediction when conflicting evidence is detected. More recently, Wang et al. (2024) propose an instruction-based framework that enables LLMs to identify conflicts, localize conflicting segments, and generate distinct responses for conflicting scenarios.
4.2 Hallucination
4.2.1 Origins
Factual Hallucinations
Factual hallucinations arise when a model’s output contradicts real-world facts. Lin et al. (2022) present TruthfulQA, an adversarial QA benchmark, and show that even top-performing models like GPT-3 were truthful on only 58% of questions, compared to 94% for humans. Pagnoni et al. (2021) construct FRANK, a dataset for identifying factual errors in summarization, while Honovich et al. (2021) extend QAGS to dialogue by leveraging question generation and entailment for factual consistency evaluation. To assess factual knowledge and reasoning in LLMs, Hu et al. (2024) introduce Pinocchio, a large benchmark covering multiple domains, timelines, and languages, revealing challenges in composition, temporal reasoning, and robustness. Mallen et al. (2023) further find that models struggle with less common factual knowledge, with retrieval augmentation significantly improving performance in such cases.
Contextual Hallucinations
Contextual hallucinations occur when generated text contradicts the given input context, such as in summarization, translation, and generation tasks. Maynez et al. (2020) find that summarization models frequently generate content unfaithful to input documents, with 64% of summaries containing unsupported information. In machine translation, Raunak et al. (2021) analyze hallucinations caused by source perturbations and training noise, and find that slight modifications to input data could trigger off-topic translations. Similarly, Dale et al. (2023) introduce HalOmi, a multilingual benchmark for hallucination and omission detection in machine translation, showing that prior hallucination detectors often fail across different language pairs. In generation tasks, Liu et al. (2022a) propose a novel token-level, reference-free hallucination detection task and dataset (HADES) for free-form text generation, and Niu et al. (2024) introduce RAGTruth, a comprehensive corpus designed for analyzing word-level hallucinations across various domains and tasks within standard Retrieval-Augmented Generation (RAG) frameworks.
4.2.2 Implications and Mitigation
Implications
Hallucinations threaten trust, safety, and the integrity of AI-powered workflows. Hallucinated outputs can rapidly spread false information. For instance, in 2023, an AI-generated image purporting to show an explosion near the Pentagon went viral, briefly causing public panic and even a stock market dip before being debunked Sun et al. (2024). Hallucinations directly degrade the performance of downstream applications like abstractive summarization. Studies have found that a large portion of generated summaries contain unsupported facts, misleading readers and propagating misinformation in news and scientific dissemination Kryscinski et al. (2020).
Mitigation
Mitigating hallucinations in language models has been approached through various strategies, including knowledge disentanglement Neeman et al. (2023), retrieval augmentation Lewis et al. (2020); Shuster et al. (2021), knowledge graphs Guan et al. (2024), and improved verification methods Kryscinski et al. (2020); Wang et al. (2020); Laban et al. (2022); Manakul et al. (2023). DisentQA enhances robustness by training models to separate internal memory from external context, improving accuracy in conflicting knowledge scenarios Neeman et al. (2023). Retrieval-Augmented Generation (RAG) mitigates factual inconsistencies by integrating external sources like Wikipedia Lewis et al. (2020) or incorporating a neural search module into chatbot responses Shuster et al. (2021). In addition, Guan et al. (2024) demonstrate how retrofitting LLM outputs using structured knowledge graphs can correct factual inconsistencies, particularly in complex reasoning tasks. For hallucination detection methods, FactCC and QAGS introduce automated methods using synthetic data and question-answer validation to assess factual consistency Kryscinski et al. (2020); Wang et al. (2020). SummaC refines entailment-based scoring Laban et al. (2022), and SelfCheckGPT detects hallucinations by sampling multiple model outputs and checking for agreement without external references Manakul et al. (2023).
5 Connections, Challenges and Directions
Given the significance and impact of conflicts in NLP, we advocate for increased attention to the development of conflict-aware and robust AI systems. In this section, we highlight specific challenges by connecting different types of conflicts and propose concrete research directions to address them.
Culturally Robust LLMs
Among the challenges outlined in this survey, the development of culturally robust LLMs remains particularly underexplored. Cultural conflicts emerge both in naturally occurring web data and human-annotated datasets, where Western-centric distributions dominate. Prior studies reveal that LLMs often reflect the values and perspectives of Western, English-speaking populations Ramaswamy et al. (2023); Mihalcea et al. (2025); Tao et al. (2024); Naous et al. (2024), with misalignments especially pronounced for underrepresented personas and culturally sensitive topics Al Kuwatly et al. (2020). Additionally, LLMs exhibit inconsistent behavior across languages Li et al. (2024a); AlKhamissi et al. (2024); Eloundou et al. (2025), revealing internal cultural conflicts. These issues are rooted in the data: both the pre-train data and benchmark datasets commonly exhibit Western-centric biases Mihalcea et al. (2025); Faisal et al. (2022), causing models to default to Western contexts and perform poorly on less-represented regions and cultures.
However, to the best of our knowledge, no effective methodology has yet been proposed to address this issue. With the emergence of culturally distinct LLMs—such as Qwen, trained largely on Chinese data Bai et al. (2023), and Vikhr, trained on Russian data Nikolich et al. (2025)—a promising direction is model fusion across culturally diverse models to achieve greater cultural balance Wan et al. (2024b); Jiang et al. (2023a). Furthermore, advances in culture-specific LLMs and synthetic data generation offer the potential to curate more culturally representative training and evaluation datasets beyond Western-centric narratives, supporting the development of culturally robust LLMs.
Building Conflict-Aware AI Systems
As outlined in this survey, various types of conflicts can arise in a model’s input, each requiring different handling depending on the task. We argue that downstream applications should not treat all conflicts uniformly; rather, responses should be tailored to the conflict type. For instance, conflicts due to ambiguity should elicit clarification questions, factual contradictions should trigger reasoning over evidence, and opinion-based disagreements should induce balanced, multi-perspective responses. Realizing such capabilities requires models to be aware of the potential conflicts and classify them according to a systematic taxonomy. Yet, current research lacks frameworks to distinguish and operationalize these conflict types. Our proposed taxonomy offers a foundational step toward enabling conflict-aware systems that can recognize, interpret, and appropriately address diverse conflicts in downstream applications.
6 Conclusion
We present a unified view of conflicting information in NLP, organizing the landscape into conflicts originating from (i) natural texts on the web, (ii) human annotations, and (iii) model interactions. This taxonomy connects lines of work that are often studied in isolation and clarifies how conflicts arise, what they imply for reliability, and how current methods aim to mitigate them. Our synthesis argues that conflict awareness should guide the full pipeline: data collection that preserves disagreement and multiple perspectives, models that detect and categorize conflicts and respond with clarification, reasoning, or balanced presentation, and evaluations that measure calibration under disagreement, robustness to contradictory evidence, and cultural coverage across languages and regions.
Limitations
Conflicting information is present both in the data that models rely on and in their generated outputs. While we strive to account for all potential conflict scenarios, some cases may inevitably be overlooked. Additionally, due to space constraints, we cannot provide an exhaustive discussion of the literature on each specific type of conflict. Instead, we adopt a broader perspective, examining various types of conflicts to identify connections, challenges, and future directions.
Acknowledgment
This research is based upon work supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via 2022-22072200003. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein. This work is also supported by NSF grant #IIS-2135581.
Appendix A Summary Tables
In the summary tables, dataset covers prior work that proposed datasets and benchmark, method covers work that focus on mitigation strategies, and analysis presents work that aim at providing insights through experiments.
Table 1: Datasets, methods, and analysis for conflicts in natural texts
Conflict Type Sub-type Category Work
Factual Ambiguity Dataset SituatedQA Zhang and Choi (2021)
AmbigQA Min et al. (2020)
Time-sensitive QA Chen et al. (2021)
StreamingQA Liska et al. (2022)
Real-time QA Kasai et al. (2023)
Method Disambiguate then answer Cole et al. (2023)
Time-aware LM Dhingra et al. (2022)
Contradictory Evidence Dataset QACC Liu et al. (2024)
Contra-Doc Li et al. (2024b)
WhoQA Pham et al. (2024)
Machine-generated Pan et al. (2023)
Machine-generated Wan et al. (2024a)
Machine-generated Hong et al. (2024)
Machine-generated Jiayang et al. (2024)
Entity-substitution Chen et al. (2022a)
Rule-based Kazemi et al. (2023)
Method Finetuned Calibrator Chen et al. (2022a)
Finetuned w/ Explanation Liu et al. (2024)
Finetuned discriminator Hong et al. (2024)
Analysis Confirmation bias Xie et al. (2024)
Opinion Perspectives Dataset PERSPECTRUM Chen et al. (2019)
Multi-OpEd Liu et al. (2021)
NeuS Lee et al. (2022)
ConflictingQA Wan et al. (2024a)
Reddit Plepi et al. (2024)
Method Multi-task learning Liu et al. (2021)
Opinion summarization Jiang et al. (2023b)
Tailored generation Plepi et al. (2024)
Framing Bias Dataset MFC Card et al. (2015)
GVFC Liu et al. (2019)
BASIL Fan et al. (2019)
Method Multifaceted analysis Milbauer et al. (2021)
Pre-training Liu et al. (2022b)
Disentanglement Chen et al. (2023)
Multi-task learning Lee et al. (2022)
Analysis Sentence-level Lei et al. (2022)
Table 2: Datasets, methods, and analysis for conflicts in human-annotated texts
Conflict Type Sub-type Category Work
Human-Annotated Disagreement Dataset Twitter Sandri et al. (2023)
RLHF Dsouza and Kovatchev (2025)
DiscoGeM Yung and Demberg (2025)
NLI Pavlick and Kwiatkowski (2019)
Method Probabilistic model Sheng et al. (2008)
Multi-task Mostafazadeh Davani et al. (2022)
Soft labels Uma et al. (2021)
Soft labels Fornaciari et al. (2021)
Analysis Survey Uma et al. (2021)
Survey Klie et al. (2023)
Offensive language Sandri et al. (2023)
NLI Jiang and de Marneffe (2022)
Task design Dsouza and Kovatchev (2025)
Free choice Yung and Demberg (2025)
Personal belief Sap et al. (2022)
Demographic data Wan et al. (2023)
Biases Dataset Gender Buolamwini and Gebru (2018)
Argument mining Thorn Jakobsen et al. (2022)
POPQUORN Pei and Jurgens (2023)
Analysis Western-centric Faisal et al. (2022)
Toxicity Wan et al. (2023)
Racist outcome Sap et al. (2019)
Table 3: Datasets, methods, and analysis for conflicts during model interactions
Conflict Type Sub-type Category Work
Knowledge Context vs. Memory Dataset Entity substitution Longpre et al. (2021)
Entity substitution Chen et al. (2022a)
Instruction-based Wang et al. (2024)
Misinformation injection Pan et al. (2023)
KRE Ying et al. (2024)
context-conflicting Tan et al. (2024)
Method Data Augmentation Longpre et al. (2021)
Abstention Chen et al. (2022a)
Instruction-based Wang et al. (2024)
Within & Across Analysis LM-vs-LM fact-checking Cohen et al. (2023)
Cross-modality Zhu et al. (2024)
Intra-model contradiction Zhao et al. (2024)
Hallucination Factual Dataset TruthfulQA Lin et al. (2022)
FRANK Pagnoni et al. (2021)
$q^{2}$ Honovich et al. (2021)
Pinocchio Hu et al. (2024)
MiniCheck Tang et al. (2024)
Method RAG Lewis et al. (2020)
RAG Shuster et al. (2021)
Knowledge graph Guan et al. (2024)
Disentanglement Neeman et al. (2023)
QA validation Kryscinski et al. (2020)
QA validation Wang et al. (2020)
Entailment-based Laban et al. (2022)
SelfCheckGPT Manakul et al. (2023)
Analysis Less popular entities Mallen et al. (2023)
Contextual Dataset HalOmi Dale et al. (2023)
HADES Liu et al. (2022a)
RAGTruth Niu et al. (2024)
Method Context-aware decoding Shi et al. (2024)
Long context Liu et al. (2025)
Context-DPO Bi et al. (2024)
CR-DPO Huang et al. (2025)
Analysis Summarization Maynez et al. (2020)
Translation Raunak et al. (2021)
Appendix B Survey Methodology
Our objective is to map the broad and diverse landscape of textual conflicts in NLP, identify central implications and gaps, and outline a research agenda. We surveyed research indexed in Google Scholar, the ACL Anthology, and OpenReview, covering peer-reviewed work in major ML and NLP venues such as ICLR and ACL, as well as recent manuscripts on arXiv. Rather than relying on predetermined keyword filters, we primarily used citation chaining: starting from influential surveys corresponding to the three sources of conflict considered in this work (natural web text, human annotations, and model interactions), we applied both backward and forward chaining using Google Scholar and venue indexes. Paper selection did not follow a rigid checklist. The two authors independently screened and included papers based on inclusion criteria that prioritized recency, citation impact, and perceived influence on the area, then reconciled disagreements through discussion. This methodology emphasizes coverage and structure over exhaustiveness and is intended to synthesize a rapidly evolving field while making the scope of inclusion explicit and surfacing open problems.